hi123654789:

Consider the following pair of equations: −x − y = −5 y = x + 1 If the two equations are graphed, at what point do the lines representing the two equations intersect?

Kylo:

well have you tried working it out if not i can help

hi123654789:

i tried, but i'm stuck still

Kylo:

well here is what ive found when you solve it you should get this i think so what do you think the answer is

hi123654789:

(3, 2) (2, 3) (−2, 3) (3, −2) those are my choices

Kylo:

theres no negative number in it so there

hi123654789:

So (2,3)?

Kylo:

yes! : )

Kylo:

glade to help you

hi123654789:

thank you!

Kylo:

no problem if you have anymore questions related to this kind of math i can help im pretty good at that kind of stuff

hi123654789:

i do have one more

Kylo:

okay im here to help if you need it

hi123654789:

Consider the following system of equations: y = −x + 2 y = 3x + 1 Which description best describes the solution to the system of equations? Lines y = −x + 2 and y = 3x + 1 intersect the x-axis. Lines y = −x + 2 and y = 3x + 1 intersect the y-axis. Line y = −x + 2 intersects line y = 3x + 1. Line y = −x + 2 intersects the origin.

Kylo:

here is what i found

Kylo:

they do not intersect at the origin

hi123654789:

Lines y = −x + 2 and y = 3x + 1 intersect the x-axis?

Kylo:

i think so! : )

Kylo:

i have to go now if you ever have any questions feel free to dm me!

hi123654789:

You're good at this. also thank you again have an amazing day

Kylo:

you too!

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hi123654789 Lines y = −x + 2 and y = 3x + 1 intersect the x-axis? \(\color{#0cbb34}{\text{End of Quote}}\) No, that is not correct Take a closer look at the graph: https://www.desmos.com/calculator/tyqrhgzcvn Does it intersect at the x-axis? y-axis? origin? or just intersect?

1 attachment